OpenStudy (anonymous):

y varies inversely as twice x. When x = 2, y = 5. Find y when x = 10. Please helppp me

OpenStudy (anonymous):

y=k/2x

OpenStudy (anonymous):

Now try

myininaya (myininaya):

y=k/(2x) not just plug in x=2 and y=5 to find the constant k

OpenStudy (anonymous):

y = 1 y = 20 y = 8 y = 10 these are the possible answers

myininaya (myininaya):

now*

OpenStudy (anonymous):

idkk

myininaya (myininaya):

what do you for k when pluggin in x=2 and y=5

OpenStudy (anonymous):

5/4

myininaya (myininaya):

you see we have y=k/(2x) just put a 5 where that y is and just put a 2 where that x is

myininaya (myininaya):

so you would have 5=k/(2*2) right? now solve for k

OpenStudy (anonymous):

5=k/4?

myininaya (myininaya):

ok solve for k

OpenStudy (anonymous):

20?

myininaya (myininaya):

right k=20 gj so we have y=20/(2x) or you can reduce this fraction y=10/x

myininaya (myininaya):

we want to find y when x=10

myininaya (myininaya):

so if x=10 then y=?

OpenStudy (anonymous):

so what do i do?

myininaya (myininaya):

plug in 10 where u see x

OpenStudy (anonymous):

so y=1?

myininaya (myininaya):

yes because 10/10=1

OpenStudy (anonymous):

k thank you so much